The manager of a bulk foods establishment sells a trail mix for $6 per pound and premium cashews for $12 per pound. The manager wishes to make a 135-pound trail mix-cashew mixture that will sell for $10 per pound. How many pounds of each should be used?
Answer by checkley77(12844)

12X+6(135-X)=135*10
12X+810-6X=1350
6X=1350-810
6X=540
X=540/6
X=90 LBS. OF $12 CASHEWS IS USED.
135-90=45 LBS. OF $6 TRAIL MIX IS USED.
PROOF:
12*90+6*45=135*10
1080+270=1350
1350=1350